# Quasinormal Modes and Dynamical Evolution of Scalar Fields in the Einstein–Bumblebee Theory with a Cosmological Constant

### v1

@@ III.1 Wenzel–Kramers–Brillouin (WKB) Approximation Method
In 1926, Wenzel, Kramers, and Brillouin introduced a semiclassical approximation method for solving the Schrödinger equation. In the field of quantum mechanics, this method is extensively used for semiclassical calculations, such as calculating the penetration rates of bound state potential barriers. The WKB approximation integrates elements of both classical and quantum theories, offering an effective tool for addressing challenges that are difficult to resolve through purely classical or quantum approaches. This is particularly advantageous in dealing with potential barrier penetration and quantum tunneling effects.
@@
III.2 Quasinormal Mode Numerical Results
@@
Figures 2(a)-2(f) demonstrate that as the Lorentz violation parameter $\ell$ increases, both the real and imaginary parts of the quasinormal modes in the lower modes tend to decrease. In higher modes, changes in the real part of the quasinormal modes are not significant, but the imaginary parts decrease rapidly. Additionally, when the cosmological constant $\Lambda$ increases, the overall behavior of the quasinormal modes tends to decrease. The impacts on the dipole and quadrupole modes are similar in trend but different in magnitude. Anomalous behavior is observed in the monopole mode, where the cosmological constant $\Lambda$ seems to have a negligible effect on the imaginary part. The validity of these results will be verified in the next section through fitting dynamical waveforms.
@@ IV Dynamical Evolution
Figures 3(a)-3(i) display the dynamical evolution waveforms of scalar field perturbations under various black hole parameter conditions. Here, we perform a coordinate transformation $(u,v)\rightarrow(t_{*},r_{*})$ on the result obtained for $\psi(u,v)$ from equation (21) to restore it to the same spacetime coordinates used during the calculation of QNMs. These waveforms correspond to cases where the cosmological constant $\Lambda$ takes values of $\Lambda=0$, $\Lambda=10^{-3}$, and $\Lambda=10^{-2}$ and where the Lorentz violation parameter, $\ell$, is set to $\ell=0$, $\ell=0.5$, and $\ell=1$, under the quadrupole mode ($L=2$). We have added the logarithmic evolution plot of the scalar perturbation wave function for the $L=2$ mode in Figures 4(a)-4(i), where the linear decay of the amplitude $|\psi(t)|$ directly corresponds to the extraction of the imaginary part of the quasi-normal mode frequency, $\omega_{I}$. When the cosmological constant $\Lambda$ is zero, the images generally exhibit significant tailing phenomena. However, for moderate values of the cosmological constant, tailing only occurs when the Lorentz violation parameter $\ell$ is small. Under conditions of a relatively large cosmological constant, the tailing phenomenon completely disappears. This finding indicates that in de Sitter spacetimes, the magnitude of the cosmological constant substantially dictates the occurrence of tailing, whereas Lorentz-symmetry breaking can make the tailing less pronounced. As the breaking factor $\ell$ increases, the decay rate of the field decreases with a fixed cosmological constant, reflecting the physical manifestation of the decrease in the imaginary part of the quasinormal mode frequencies.
@@
Table 1: A comparison between the WKB approximation and waveform fitting for the $L=2$ mode under different cosmological constants and Lorentz violation parameters.
Cosmological Constant    Breaking Factor       WKB Approximation       Fitting Results
$\Lambda=0$ $\ell=0$ $0.483642-0.096761\,i$ $0.482891-0.0966647\,i$
$\ell=0.5$ $0.479592-0.078854\,i$ $0.479536-0.078892\,i$
$\ell=1$ $0.477556-0.0682362\,i$ $0.475290-0.0683287\,i$
$\Lambda=10^{-3}$ $\ell=0$ $0.481284-0.0963911\,i$ $0.481282-0.0964598\,i$
$\ell=0.5$ $0.477306-0.0785431\,i$ $0.477337-0.0784671\,i$
$\ell=1$ $0.473508-0.0679514\,i$ $0.473536-0.0680987\,i$
$\Lambda=10^{-2}$ $\ell=0$ $0.459636-0.0928598\,i$ $0.459627-0.0928873\,i$
$\ell=0.5$ $0.456132-0.0755409\,i$ $0.456325-0.0756194\,i$
$\ell=1$ $0.454652-0.0652961\,i$ $0.454691-0.0654512\,i$
Finally, the complex frequencies of these waveforms are determined, and the results are compared with those from the WKB approximation. Here, a modified exponential decay function is used to fit the data:
$Q(t)=e^{\sqrt{1+\ell}\omega_{I}t}A_{l}\sin\left(\sqrt{1+\ell}\omega_{R}t+B_{l}\right).$ (23)
@@
where $\omega_{R}$ and $\omega_{I}$ represent the real and imaginary components of the quasinormal mode frequencies, respectively. Last, we present the fitted results and compare them with those obtained using the WKB method, as shown in Table 1. Despite the potential for inherent errors in the numerical computation process and the limited number of parameters in the fitting procedure, the results obtained via the WKB approximation and the finite difference method meet the required accuracy. Furthermore, our fitting results demonstrate the consistency between the two methods.
@@ V Conclusion
In this work, we investigated scalar field perturbations within the Einstein–Bumblebee theory by incorporating a cosmological constant. We derived the equations of motion for the scalar field and solved for the eigenfrequencies via both the WKB approximation and the finite difference method. Moreover, we systematically analyzed the dependence of quasinormal mode frequencies on black hole parameters and numerically simulated dynamic waveforms characterizing scalar perturbation evolution in this spacetime. Furthermore, the research methods and results of this paper could be extended to other black hole spacetimes within this theoretical framework, particularly to rotating black holes. Given the presence of superradiance phenomena in rotating black holes Ding et al. (2020), their dynamic stability remains a significant issue in black hole physics, which needs research and discussion in the future.

### v2

@@ III.2 Wenzel—Kramers—Brillouin (WKB) Approximation Method
Alternatively, to verify the results of the matrix method, we can employ the WKB method to calculate the quasi-normal modes when $\mu=0$. In 1926, Wenzel, Kramers, and Brillouin introduced a semiclassical approximation method for solving the Schrödinger equation. In the field of quantum mechanics, this method is extensively used for semiclassical calculations, such as calculating the penetration rates of bound state potential barriers. The WKB approximation integrates elements of both classical and quantum theories, offering an effective tool for addressing challenges that are difficult to resolve through purely classical or quantum approaches. This is particularly advantageous in dealing with potential barrier penetration and quantum tunneling effects.
@@
III.3 Quasinormal Mode Numerical Results
@@
Figs. 2(a)—2(f) show the QNMs of a massless scalar field using the WKB method. As the Lorentz violation parameter $\ell$ increases, both the real and imaginary parts of the quasinormal modes in the monopole modes tend to decrease. In dipole and quadrupole modes, changes in the real part of the quasinormal modes are not significant, but the imaginary parts decrease rapidly. Additionally, when the cosmological constant $\Lambda$ increases, the overall behavior of the quasinormal modes tends to decrease. The impacts on the dipole and quadrupole modes are similar in trend but different in magnitude. Anomalous behavior is observed in the monopole mode, where the cosmological constant $\Lambda$ seems to have a negligible effect on the imaginary part. The validity of these results will be verified in the next section through fitting dynamical waveforms.
Figure 3: Complex massive scalar frequencies for the $n=0$ mode for varying values of $\mu M$.
Figs. 3(a)—3(f) show the variation of the real and imaginary parts of QNMs with increasing field mass under different Lorentz-violating parameters in the presence of a cosmological constant using the MM method with $N=30$. For the dipole and quadrupole modes, the effect of the field mass on the frequency is monotonic, with the real part increasing and the imaginary part decreasing as the field mass increases. It is noteworthy that all parameters discussed in this paper reduce the imaginary part of QNMs; however, the Lorentz-violating parameters result in a slower decrease, as observed in the figures, while the field mass leads to a more rapid decrease. For the monopole mode, the introduction of the field mass leads to significantly different results. Within the range where the field mass satisfies $\mu M<0.1$, it appears to have no effect on QNMs, whereas in the large mass regime, its influence resembles that of other modes. This is because the field mass alters the boundary behavior of the scalar field’s effective potential at infinity, which is consistent with the effect of the cosmological constant. The monopole mode, lacking angular quantum numbers, is governed by the cosmological constant in terms of the effective potential’s behavior, which accounts for the aforementioned phenomenon.
@@ IV Dynamical Evolution
Figs. 4(a)-4(i) display the dynamical evolution waveforms of scalar field perturbations under various black hole parameter conditions. Here, we perform a coordinate transformation $(u,v)\rightarrow(t_{*},r_{*})$ on the result obtained for $\psi(u,v)$ from equation (27) to restore it to the same spacetime coordinates used during the calculation of QNMs. These waveforms correspond to cases where the cosmological constant $\Lambda$ takes values of $\Lambda=0$, $\Lambda=10^{-3}$, and $\Lambda=10^{-2}$ and where the Lorentz violation parameter, $\ell$, is set to $\ell=0$, $\ell=0.5$, and $\ell=1$, under the quadrupole mode ($L=2$). We have added the logarithmic evolution plot of the scalar perturbation wave function for the $L=2$ mode in Figures 5(a)-5(i), where the linear decay of the amplitude $|\psi(t)|$ directly corresponds to the extraction of the imaginary part of the QNMs, $\omega_{I}$. When the cosmological constant $\Lambda$ is zero, the images generally exhibit significant tailing phenomena. However, for moderate values of the cosmological constant, tailing only occurs when the Lorentz violation parameter $\ell$ is small. Under conditions of a relatively large cosmological constant, the tailing phenomenon completely disappears. This finding indicates that in de Sitter spacetimes, the magnitude of the cosmological constant substantially dictates the occurrence of tailing, whereas Lorentz-symmetry breaking can make the tailing less pronounced. As the Lorentz violation parameter $\ell$ increases, the decay rate of the field decreases with a fixed cosmological constant, reflecting the physical manifestation of the decrease in the imaginary part of the quasinormal mode frequencies.
@@
Table 1: A comparison is conducted between the matrix method, the WKB approximation, and the waveform fitting for the $L=2$ mode with different values of the cosmological constant and Lorentz-violating parameters.
      Matrix Method       WKB       Fitting
$\Lambda=0$ $\ell=0${}~{}~{}~{}~{}~{}~{}~{}-~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}- $0.483642-0.096761\,i$ $0.482891-0.096664\,i$
$\ell=0.5${}~{}~{}~{}~{}~{}~{}~{}-~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}- $0.479592-0.078854\,i$ $0.479536-0.078892\,i$
$\ell=1${}~{}~{}~{}~{}~{}~{}~{}-~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}- $0.477556-0.068236\,i$ $0.475290-0.068328\,i$
$\Lambda=10^{-3}$ $\ell=0$ $0.482488-0.097261\,i$ $0.481284-0.096391\,i$ $0.481282-0.096459\,i$
$\ell=0.5$ $0.478058-0.080636\,i$ $0.477306-0.078543\,i$ $0.477337-0.078467\,i$
$\ell=1$ $0.474743-0.071302\,i$ $0.473508-0.067951\,i$ $0.473536-0.068098\,i$
$\Lambda=10^{-2}$ $\ell=0$ $0.459638-0.092853\,i$ $0.459636-0.092859\,i$ $0.459627-0.092887\,i$
$\ell=0.5$ $0.456312-0.075539\,i$ $0.456132-0.075540\,i$ $0.456325-0.075619\,i$
$\ell=1$ $0.454652-0.065296\,i$ $0.454652-0.065296\,i$ $0.454691-0.065451\,i$
@@
Finally, we present the fitted results and compare them with those obtained using the Matrix Method and the WKB Approximation, as shown in Table 1. Due to fundamental differences in the boundary conditions described earlier, when $\Lambda=0$, the results of the Einstein-Bumblebee spacetime cannot be directly derived as the limiting case of the Einstein-Bumblebee dS spacetime; therefore, in the Matrix Method adopted in this paper, we cannot obtain results for the $\Lambda=0$ case. When the cosmological constant is small, the accuracy of the Matrix Method for $N=30$ is poor, but this low accuracy can be improved by increasing the matrix order. However, when the cosmological constant is large, the Matrix Method exhibits very high accuracy, even with a low matrix order such as $N=15$. Despite the potential for inherent errors in the numerical computation process and the limited number of parameters in the fitting procedure, the results obtained via the WKB Approximation and the Finite Difference Method meet the required accuracy. Furthermore, our fitting results demonstrate consistency between the WKB Approximation and the Finite Difference Method.
@@ V Conclusion
In this work, we investigated scalar field perturbations within the Einstein—Bumblebee theory by incorporating a cosmological constant. We derived the equation of motion for a massive scalar field, solved for the eigenfrequencies in the massive case using the matrix method, and verified the results in the massless case using the WKB approximation and finite difference method. Moreover, we systematically analyzed the dependence of quasinormal mode frequencies on black hole parameters and numerically simulated dynamic waveforms characterizing scalar perturbation evolution in this spacetime. Furthermore, the research methods and results of this paper could be extended to other black hole spacetimes within this theoretical framework, particularly to rotating black holes. Given the presence of superradiance phenomena in rotating black holes Ding et al. (2020), their dynamic stability remains a significant issue in black hole physics, which needs research and discussion in the future.